# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ VI-A2 The Projected Subgradient Algorithm
Problem $\widetilde{{\cal S}}_{o}^{\prime}$ is a non-convex-concave min-max problem, whose structure is the same as $\widetilde{{\cal S}}_{\bf w}^{\prime}$ in (16). Thus, we can apply PSA with the updating steps similar to (17) and (18). In particular, the updating procedure at iteration $j$ is given by
${\bf y}^{(j)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}g({\bf x}^{(j)},{\bf y});$ (34)
${\bf x}^{(j+1)}=\Pi_{{\cal U}}\big{(}{\bf x}^{(j)}-\eta\nabla_{{\bf x}}g({\bf x}^{(j)},{\bf y}^{(j)})\big{)}$ (35)
@@
where $\eta>0$ is the step size, and $\Pi_{{\cal U}}({\bf x})$ is the projection of ${\bf x}$ onto the feasible set ${\cal U}$. Since the constraints on ${\bf a}$ and ${\bf e}$ in ${\cal U}$ are separate, $\Pi_{{\cal U}}({\bf x})$ can be performed for ${\bf a}$ and ${\bf e}$ separately:
$\Pi_{{\cal U}}({\bf x})=\begin{cases}{\bf a}&\mbox{if }\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\\ \sqrt{\frac{P}{P_{\text{tot}}}}{\bf a}&\text{o.w.}\\ e_{m}&\mbox{if }|e_{m}|\leq 1,\ m\in{\cal M}\\ \frac{e_{m}}{|e_{m}|}&\text{o.w.}\end{cases}.$ (36)
@@
$g({\bf x},{\bf y}^{(j)})=\varphi_{\hat{i}\hat{k}}({\bf x})=\max_{i,k}\varphi_{ik}({\bf x}),$ (37)
for some $(\hat{i},\hat{k})$. Then, the gradient $\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ is given by
$\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ $=\nabla_{{\bf x}}\varphi_{\hat{i}\hat{k}}({\bf x})=\Bigl{[}\nabla_{{\bf a}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\!,\nabla_{{\bf e}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\Bigl{]}^{H}.$ (38)
For $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$, note that the expression for $\varphi_{ik}({\bf x})$ is the same as $\phi_{ik}({\bf e})$. Thus, the gradient $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ can be computed using (IV-B2), except that ${\bf w}_{i}$ is computed using (29), and $\zeta$ is replaced by $\delta$.
To compute $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$, we rewrite $\varphi_{ik}({\bf x})$ in (32) as
$\varphi_{ik}({\bf x})=$ $-\frac{1}{\gamma_{ik}}\frac{{\bf a}_{i}^{H}\widetilde{{\bf A}}_{iik}{\bf a}_{i}}{\sum_{j\neq i}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{jik}{\bf a}_{j}+\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$
where $\widetilde{{\bf A}}_{jik}\triangleq\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}^{H}\widetilde{{\bf C}}_{j}$, with $\tilde{{\bf h}}_{ik}={\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e}$ being the equivalent channel. Then, we have
$\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ $=[\nabla_{{\bf a}_{1}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x}),\cdots,\nabla_{{\bf a}_{G}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})]^{H}$ (39)
@@
$\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})\!=\!\begin{cases}\displaystyle-\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}}}{\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}}&\mbox{if }i=\hat{i}\\[15.00002pt] \displaystyle\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{({\bf a}_{\hat{i}}^{H}\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}})\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}}{\bigl{(}\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}\bigl{)}^{2}}&\text{o.w.}\end{cases}$ (40)
From the above, we note that the updates in (34) and (35) can be computed using closed-form evaluations.
VI-A3 Final Processing
Since we are solving $\widetilde{{\cal S}}_{o}$, which is a relaxed problem of ${\cal S}_{o}$, the computed solution ${\bf e}^{\star}$ may not satisfy (26c) for ${\cal S}_{o}$. Thus, we require this final processing step to obtain a feasible solution to ${\cal S}_{o}$. Let ${\bf x}^{\star}=[{{\bf a}^{\star}}^{H},{{\bf e}^{\star}}^{H}]^{H}$ be the solution produced by the updating procedure in (34) and (35). If ${\bf e}^{\star}$ does not satisfy (26c), we have the following step to obtain a feasible solution based on $({\bf a}^{\star},{\bf e}^{\star})$:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which takes the phase component of each element in ${\bf e}^{\star}$;
  • Given ${\bf e}^{\text{final}}$, we solve ${\cal S}_{o}^{\text{eq}}$ in (VI-A) w.r.t. ${\bf a}$ and obtain the final beamforming solution ${\bf w}^{\text{final}}$. Note that in this case, ${\cal S}_{o}^{\text{eq}}$ is reduced to the following conventional MMF problem:
    ${\cal S}^{\text{eq}}({\bf e}^{\text{final}}):\max_{{\bf a}}$ $\min_{i,k}\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (41)
    where $\textrm{SINR}_{ik}$ is given in (12) with the equivalent channel $\tilde{{\bf h}}_{ik}$ computed based on ${\bf e}^{\text{final}}$. As mentioned earlier, the PSA-based fast algorithm for this MMF problem is provided in [12]. We directly apply this algorithm to compute the solution ${\bf a}^{\text{final}}$ to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$. For initialization in this algorithm, we use ${\bf a}^{\star}$ as a good initial point to this algorithm to achieve fast convergence and good performance.
We summarize our proposed fast algorithm for the MMF problem ${\cal S}_{o}$ in Algorithm 2. Again, based on the existing convergence analysis [12], the PSA updating procedure in (34) and (35) is guaranteed to converge in finite time to a point at the vicinity of a stationary point for $\widetilde{{\cal S}}_{o}$.
Algorithm 2 The Proposed Fast Algorithm for RIS-Assisted MMF Problem ${\cal S}_{o}$.
1:Initialization: Set feasible point ${\bf x}^{(0)}$; Set $n=0$.
@@ VI-A3 Final Processing
3:     Update ${\bf x}^{(n)}$ using (34) and (35) to obtain ${\bf x}^{(n+1)}$.
4:     Set $n\leftarrow n+1$.
5:until convergence
6:Set ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{(n)})$.
7:Solving ${\cal S}({\bf e}^{\text{final}})$ to obtain ${\bf w}^{\text{final}}$.
8:return (${\bf w}^{\text{final}}$, ${\bf e}^{\text{final}}$).
VI-B Discussions
VI-B1 Initialization
@@
The proposed updating procedure in (34) and (35) requires an initial point ${\bf x}^{(0)}=[{{\bf a}^{(0)}}^{H},{{\bf e}^{(0)}}^{H}]^{H}$. This initial point can be generated randomly, as the projection (35) will ensure that the subsequent points $\{{\bf x}^{(n)}\}$ are feasible. However, a good initial point improves the rate of convergence. Thus, instead of a random initial point, we consider the following approach.
We first generate ${\bf e}^{(0)}$ with a random phase for each element. Given ${\bf e}^{(0)}$, ${\cal S}_{o}^{\text{eq}}$ in (VI-A) is reduced to the MMF problem w.r.t. ${\bf a}$, which is similar to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$ in (41). It can be further equivalently written as
${\cal S}^{\text{eq}}({\bf e}^{(0)}):$ $\max_{{\bf a},t}\ t$ s.t. $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq t,\ k\in{\cal K}_{i},i\in{\cal G}$ (42a)
$\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (42b)
@@
Note that any feasible solution to the above problem is feasible to $\widetilde{{\cal S}}_{o}$. Thus, we generate ${\bf a}^{(0)}$ in an attempt to solve this MMF problem, which is to solve its inverse problem, i.e., the QoS problem similar to ${\cal P}_{\bf e}$ in (III-B1) for given $t$, along with bi-section search over $t$ [28, 11]. However, since our goal is only to generate a feasible point ${\bf a}^{(0)}$ for $\widetilde{{\cal S}}_{o}$, we only need to solve the QoS problem [^2] To solve the QoS problem, we use the classical approach of SDR with the Gaussian randomization procedure. along with one round of bi-section search over $t$ to obtain the initial point ${\bf a}^{(0)}$.
VI-B2 Computational Complexity
@@
For the updates in (34) and (35) at each iteration, the update ${\bf y}^{(j)}$ is obtained in (37) by finding the maximum $\varphi_{ik}({\bf e})$ among $K_{\text{tot}}$ users, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+2N+4)$ flops. The update in (35) requires to calculate $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ and $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$. Note that $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ is computed using (IV-B2), which requires $2GM+2M$ flops as discussed in Section IV-C2.ii). To compute $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ in (40), we need ${\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $j\in{\cal G}$, which is part of the SINR expression that has already been computed in $\varphi_{ik}({\bf e})$. We also need to calculate $\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}=\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}\tilde{{\bf h}}_{\hat{i}\hat{k}}^{H}\widetilde{{\bf C}}_{i}{\bf a}_{i}$, for $i\in{\cal G}$. Note that $\widetilde{{\bf C}}_{i}$’s only need to be computed once before the update procedure, and $\tilde{{\bf h}}_{\hat{i}\hat{k}}$ is already computed as part of the SINR expression. Thus, compute $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}$’s for $i\in{\cal G}$ require $K_{\text{tot}}(N+1)$. Lastly, we need $K_{\text{tot}}+G$ flops to calculate the final value of $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ for all ${\bf a}_{i}$’s. In summary, the complexity in obtaining the update in (35) requires $2(G+1)M+K_{\text{tot}}(N+2)+G$ flops. Thus, the overall leading complexity of the updating procedure in each PSA iteration is dominated by computing ${\bf y}^{(j)}$ in (37), which is $GK_{\text{tot}}N+K_{\text{tot}}MN$ flops. It grows linearly with $M$ and $N$.
VI-C General RIS-Assisted Multi-User Downlink Beamforming
Similar to the discussion in Section V about the generalization to downlink RIS-assisted multi-user beamforming for the QoS power minimization problem, our proposed fast algorithm (Algorithm 2) is directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the max-min fair objective for both unicast and multicast scenarios. Specific to the unicast case with $K_{i}=1,\forall i$, the computational complexity per iteration in the proposed iterative algorithm is reduced to $GMN+G^{2}N$ flops, which is computationally cheap.
For the multi-user unicast scenario, the MMF problem for RIS-assisted downlink beamforming under various system settings has been considered in the recent literature, and various approaches have been proposed [22, 23, 24]. In [22], for the same setting considered in this work, an iterative algorithm that provides an alternating descent path is proposed to maximize the minimum SINR, and the computational complexity in each iteration as $O((GN)^{3}+M^{4})$. In comparison, the computational complexity of our proposed fast algorithm only grows linearly with $G$, $M$, and $N$. A different system model for RIS-assisted transmission is considered in [23], where only the RIS path is available and the channel correlation matrix between RIS and users are assumed to be the same. Using an approximation of the optimal BS transmit beamformer, the authors propose an algorithm based on projected gradient ascent to obtain the RIS beamformer based on the large-scale channel statistics. The computational complexity of the proposed algorithm is in the higher polynomial order of $N$ and $M$. RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where the RIS can adjust both reflecting phases and amplitudes. The AO approach is adopted, with two subproblems being solved by the second-order cone programming (SOCP) and SDR. An inexact AO approach with SCA is further considered to reduce the computational complexity. However, since these proposed approaches are second-order algorithms, their computational complexity is still high.
VII Simulation Results

### v3

@@ V-A2 Alternating Projected Subgradient Algorithm
Similar to the updating steps in (18) and (19) to solve $\widetilde{{\cal S}}_{\bf w}^{\prime}$ in (17), we can apply PSA to update ${\bf a}$ and ${\bf e}$ alternatingly. In particular, the updating procedure at iteration $n$ is given by
${\bf y}^{(n)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n)},{\bf y});$ (35)
${\bf e}^{(n+1)}=\Pi_{{\cal E}}\big{(}{\bf e}^{(n)}-\tilde{\alpha}\nabla_{{\bf e}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n)},{\bf y}^{(n)})\big{)}$ (36)
$\tilde{{\bf y}}^{(n)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n+1)},{\bf y});$ (37)
${\bf a}^{(n+1)}=\Pi_{{\cal A}}\big{(}{\bf a}^{(n)}-\tilde{\alpha}\nabla_{{\bf a}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n+1)},\tilde{{\bf y}}^{(n)})\big{)}$ (38)
@@
where $\tilde{\alpha}>0$ is the step size, $\Pi_{{\cal E}}({\bf e})$ is the projection of ${\bf e}$ onto ${\cal E}$ as in (20), and $\Pi_{{\cal A}}({\bf x})$ is the projection of ${\bf a}$ onto the feasible set ${\cal A}$, given by
$\Pi_{{\cal A}}({\bf x})=\begin{cases}{\bf a}&\mbox{if }\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\\ \sqrt{\frac{P}{P_{\text{tot}}}}{\bf a}&\text{o.w.}\end{cases}.$ (39)
@@
$\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})=\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})=\max_{i,k}\varphi_{ik}({\bf a},{\bf e}),$ (40)
for some pair $(\hat{i},\hat{k})$. Thus, the gradient $\nabla_{{\bf x}}\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})$, for ${\bf x}={\bf a}$ or ${\bf e}$, is given by
$\nabla_{{\bf x}}\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})$ $=\!\nabla_{{\bf x}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e}),\quad\text{for~{}}{\bf x}={\bf a}\text{~{}or~{}}{\bf e}.$ (41)
For $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$, note that the expression for $\varphi_{ik}({\bf x})$ is the same as $\phi_{ik}({\bf e})$. Thus, the gradient $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ can be computed using (IV-B2), where ${\bf w}_{i}$ is computed using (30), and $\rho$ is replaced by $\tilde{\rho}$.
To compute $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})$, we rewrite $\varphi_{ik}({\bf a},{\bf e})$ in (33) as
$\varphi_{ik}({\bf a},{\bf e})=$ $-\frac{1}{\gamma_{ik}}\frac{{\bf a}_{i}^{H}\widetilde{{\bf A}}_{iik}{\bf a}_{i}}{\sum_{j\neq i}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{jik}{\bf a}_{j}+\sigma^{2}}-\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M},$
where $\widetilde{{\bf A}}_{jik}\triangleq\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}^{H}\widetilde{{\bf C}}_{j}$. Then, we have
$\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})$ $=[\nabla_{{\bf a}_{1}}^{H}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e}),\cdots,\nabla_{{\bf a}_{G}}^{H}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})]^{H}$ (42)
@@
$\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})\!=\!\begin{cases}\displaystyle-\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}}}{\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}}&\mbox{if }i=\hat{i}\\[15.00002pt] \displaystyle\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{({\bf a}_{\hat{i}}^{H}\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}})\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}}{\bigl{(}\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}\bigl{)}^{2}}&\text{o.w.}\end{cases}$ (43)
Thus, the updates in (35)–(38) can be computed using closed-form evaluations from (39)–(43).
V-A3 Final Processing
Recall that $\widetilde{{\cal S}}_{o}$ is a relaxed problem of ${\cal S}_{o}$. Thus, the computed solution ${\bf e}^{\star}$ may not satisfy (26c) for ${\cal S}_{o}$. We use this final processing step to obtain a feasible solution to ${\cal S}_{o}$. Let $({{\bf a}^{\star}},{{\bf e}^{\star}})$ be the solution produced by the PSA procedure in (35)–(38). If ${\bf e}^{\star}$ does not satisfy (26c), we have the following step to obtain a feasible solution based on $({\bf a}^{\star},{\bf e}^{\star})$:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which obtains the phase component of each element in ${\bf e}^{\star}$;
  • Given ${\bf e}^{\text{final}}$, we solve ${\cal S}_{o}^{\text{eq}}$ in (31) w.r.t. ${\bf a}$ to obtain the final weight vector ${\bf a}^{\text{final}}$:
    ${\cal S}^{\text{eq}}({\bf e}^{\text{final}}):$ $\max_{{\bf a}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$
[…]
    where $\textrm{SINR}_{ik}$ in (1) is computed based on ${\bf e}^{\text{final}}$. This is the conventional MMF problem, which can by solved by the PSA-based fast algorithm [21]. Note that ${\bf a}^{\star}$ satisfies the power constraint in (V-A3) and thus serves as a good initial point to this algorithm to achieve fast convergence and good performance. The final beamforming solution ${\bf w}^{\text{final}}$ is computed using (30) based on ${\bf a}^{\text{final}}$.
We summarize our proposed fast algorithm for the MMF problem ${\cal S}_{o}$ in Algorithm 2. Again, similar to that discussed in Section IV-B3, we track the best solution for $({\bf a},{\bf e})$ with the minimum objective value of $\widetilde{{\cal S}}_{o}^{\prime}$ obtained so far in $n$ iterations: $\tilde{g}_{\min}^{(n)}=\min\{\tilde{g}^{(1)},...,\tilde{g}^{(n)}\}$, where $\tilde{g}^{(n)}=\max_{{\bf y}\in{\cal Y}}f({\bf a}^{(n)},{\bf e}^{(n)},{\bf y})$. Since, $\tilde{g}_{\min}^{(j)}$ is monotonically decreasing and is bounded below, we terminate the algorithm when $\tilde{g}_{\min}^{(j)}$ converges.
Algorithm 2 The APSA Algorithm for RIS-Assisted MMF Problem ${\cal S}_{o}$.
1:Initialization: Set feasible point $({\bf a}^{(0)},{\bf e}^{(0)})$; Set $\tilde{\rho}$ and $\tilde{\alpha}$; Set $n=0$.
@@ V-A3 Final Processing
3:     Compute $\varphi_{ik}({\bf a}^{(n)},{\bf e}^{(n)})$, $k\in{\cal K}_{i}$, $i\in{\cal G}$. Obtain pair $(\hat{i},\hat{k})$ as in (40).
4:     Update ${\bf e}^{(n+1)}$ using (IV-B2) and (20).
5:     Compute $\varphi_{ik}({\bf a}^{(n)},{\bf e}^{(n+1)})$, $k\in{\cal K}_{i}$, $i\in{\cal G}$. Obtain pair $(\hat{i},\hat{k})$ as in (40).
6:     Update ${\bf a}^{(n+1)}$ using (43) and (39).
7:     Set $n\leftarrow n+1$.
8:until convergence
9:Set ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{(n)})$.
10:Solve ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$ to obtain ${\bf a}^{\text{final}}$.
11:Compute ${\bf w}^{\text{final}}$ using (30).
12:return (${\bf w}^{\text{final}}$, ${\bf e}^{\text{final}}$).
V-B Algorithm Discussion
V-B1 Initialization
@@
The proposed updating procedure requires an initial point $({{\bf a}^{(0)}},{{\bf e}^{(0)}})$. This initial point can be generated randomly, as the projections $\Pi_{{\cal E}}$ and $\Pi_{Ac}$ will ensure that the subsequent points $\{({{\bf a}^{(n)}},{{\bf e}^{(n)}})\}$ are feasible. However, a good initial point help accelerate the convergence. Thus, instead of a random initial point, we consider the following approach.
@@
We first generate ${\bf e}^{(0)}$ with a random phase for each element. Given ${\bf e}^{(0)}$, ${\cal S}_{o}^{\text{eq}}$ in (31) is reduced to the MMF problem w.r.t. ${\bf a}$, i.e., ${\cal S}^{\text{eq}}({\bf e}^{(0)})$ as in (V-A3) by replacing ${\bf e}^{\text{final}}$ with ${\bf e}^{0}$. Note that any feasible solution ${\bf a}$ to ${\cal S}^{\text{eq}}({\bf e}^{(0)})$ is feasible to $\widetilde{{\cal S}}_{o}$. Thus, we only need to find a feasible solution. To do so, we solve the inverse problem of ${\cal S}^{\text{eq}}({\bf e}^{(0)})$, i.e., the QoS problem similar to ${\cal P}_{\bf e}$ in (6) for a given minimum SINR target $t$, along with one round bi-section search over $t$, to find the transmit power objective of the QoS problem [^3] To solve the QoS problem, we can use the classical approach of SDR with the Gaussian randomization procedure. that is less than $P$ and obtain the initial point ${\bf a}^{(0)}$.
V-B2 Computational Complexity
@@
For the updates in (35)–(38) at each iteration, the update ${\bf y}^{(j)}$ is obtained in (40) by finding the maximum $\varphi_{ik}({\bf e})$ among $K_{\text{tot}}$ users, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+2N+4)$ flops. The updates in (36) and (38) require to calculate $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ and $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$. Note that $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ is computed using (IV-B2), which requires $2GM+2M$ flops as discussed in Section IV-C2.ii). To compute $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ in (43), we need ${\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $j\in{\cal G}$, which is part of the SINR expression that has already been computed in $\varphi_{ik}({\bf e})$. We also need to calculate $\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}=\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}\tilde{{\bf h}}_{\hat{i}\hat{k}}^{H}\widetilde{{\bf C}}_{i}{\bf a}_{i}$, for $i\in{\cal G}$. Note that $\widetilde{{\bf C}}_{i}$’s only need to be computed once before the update procedure, and $\tilde{{\bf h}}_{\hat{i}\hat{k}}$ is already computed as part of the SINR expression. Thus, compute $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}$’s for $i\in{\cal G}$ require $K_{\text{tot}}(N+1)$. Lastly, we need $K_{\text{tot}}+G$ flops to calculate the final value of $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ for all ${\bf a}_{i}$’s. In summary, the complexity in obtaining the updates in (36) and (38) require $2(G+1)M+K_{\text{tot}}(N+2)+G$ flops. Thus, the overall leading complexity of the updating procedure in each PSA iteration is dominated by computing ${\bf y}^{(j)}$ in (40), which is $GK_{\text{tot}}N+K_{\text{tot}}MN$ flops. It grows linearly with $M$ and $N$.
V-B3 Special Case $K_{i}=1$
When $K_{i}=1$, $\forall i$, problem ${\cal S}_{o}$ reduces to the RIS-assisted multiuser unicast scenario for $G$ users with the max-min fair objective. In this case, the computational complexity per iteration in Algorithm 2 reduces to $GMN+G^{2}N$ flops, which grows linearly with $M$ and $N$.
V-B4 Improved Initialization Method for QoS Problem ${\cal P}_{o}$
We can adopt the APSA algorithm as an improved initialization method to AMBF (Algorithm 1) for the QoS problem. Specifically, following the random phase initialization in Section IV-C1, we have the initial point $({\bf e}^{0},{\bf w}^{0})$. We feed this point to APSA with power budget set to $P=\sum_{i=1}^{G}\|{\bf w}_{i}^{0}\|^{2}$. The output ${\bf e}^{\text{final}}$ from APSA is then used as the updated initial point for ${\bf e}$ in Algorithm 1: ${\bf e}^{0}\leftarrow{\bf e}^{\text{final}}$. Also, as the purpose of APSA here is for providing an initial point, we can set a pre-set number of iterations for Algorithm 2 for the output before waiting for the algorithm converges,
VI Simulation Results